OpenStudy (anonymous):

x^2+8x+13 factor trinomial completely

OpenStudy (anonymous):

@BrennaM

OpenStudy (campbell_st):

there is a problem in that it would only factorise if you used the general quadratic formula to find the factors...

OpenStudy (campbell_st):

this is because 13 is prime, and only has factors of 1 and 13...

OpenStudy (anonymous):

so is the polynomial prime?

OpenStudy (campbell_st):

I'd say so...

OpenStudy (anonymous):

how about x^2-4x+3

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

do you know the factors of 3 that add to - 4, (hint: they are both negative)

OpenStudy (anonymous):

-2?

OpenStudy (campbell_st):

the factors are -1 and -3 -1*-3 = 3 -1 + -3 = -4 so the factored form is (x -1)(x - 3)

OpenStudy (anonymous):

okay and the next step?

OpenStudy (anonymous):

oh thats the final answer?

OpenStudy (campbell_st):

that sure is